14. In​ 1975, interest rates were 7.85% and the rate of inflation was 12.3% in the United States. What was the real interest rate in​ 1975? How would the purchasing power of your savings have changed over the​ year? ​(Note: Be careful not to round any intermediate steps less than six decimal​ places.)
Answer #1

I have answered the question below.

Please upvote for the same and thanks!!!

Do reach out in the comments for any queries

Answer:

Real rate of interest in 1975 can be calculated as = (1.0785/1.123) - 1

=> Real rate of interest in 1975 = -3.9626001%

Also , the purchasing power would have decreased by -3.9626001% due to very high inflation
Answer #2
rr = (r - i) / (1 + i) = (7.85% - 12.3%) / 1.123 = -3.96% The purchasing power of your savings declined by 3.96% over the year. Real rate of interest in 1975 was 3.96%
source: My teaching material
answered by: anonymous